# Beyond Winning Strategies: Admissible and Admissible Winning Strategies for Quantitative Reachability Games

### v1

@@ A.16 Proof of Theorem 8
As $\operatorname{sVal}(v)=1$ only in the winning region, we have the following equivalent statement $\neg(v\in V_{win})\vee(v^{\prime}\in V_{win})$ where $v:=\operatorname{last}(h)$ and $v^{\prime}:=\delta(\operatorname{last}(h),\sigma(h))$. Thus, modifing the admissibility checking criteria in Line 9 of Alg. 1 to $\operatorname{cVal}(v^{\prime})<\min\{\operatorname{aValues}\}$ and $\neg(v\in V_{win})\vee(v^{\prime}\in V_{win})$ is sufficient to capture the admissibility criteria for $\operatorname{\mathbf{mSC}}$. Thus, from Thms. 5 and 10, we have that Alg. 4 returns the set of all admissible winning strategies, is sound and correct. ∎
A.17 Note on Alg. 2
Alg. 2 is the Value Iteration algorithm for computing optimal $\operatorname{Val}$s and optimal strategy from every state in $\mathcal{G}$ (Brihaye et al. 2017)[Adapted and Modified from Algo. 1]. Line 2 and Line 2 correspond to computation when the Env player is playing adversarially (maximally) and cooperatively (minimally), respectively. When Env is playing adversarially, Alg. 2 returns optimal $\operatorname{aVal}$ and worst-case optimal strategy $\operatorname{\mathbf{WCO}}$. For $\operatorname{aVal}$ computation, $\sigma_{win}$ is the optimal winning strategy from every state in $V_{win}$.
@@
A.18 Note on Alg. 3 and Alg. 4
Alg. 3 and Alg. 4 compute admissible and admissible winning strategies, respectively. The key components are unrolling the graph, running value iteration, and an algorithm for Depth-First Search (DFS). The DFS algorithm searches over $\mathcal{G}^{\prime}$ iteratively rather than recursively in order to avoid Python’s low recursion limit. Specifically, the algorithm explores states in preorder fashion. Alg. 2 gives the outline to compute optimal $\operatorname{aVal}$ and $\operatorname{cVal}$ and the corresponding witness strategies. See A.17 for more details. After we unroll the graph, we compute $\operatorname{aVal}$, $\operatorname{cVal}$, and $\operatorname{acVal}$ for every state and finally run the DFS algorithm for strategy synthesis.
@@
Appendix B Note on $\operatorname{\mathbf{SCO}}$ vs $\operatorname{\mathbf{SC}}$
Brenguier et al. define strongly-cooperative optimal strategies in order to show that every $\operatorname{\mathbf{SCO}}$ is admissible. Then, they proceed to shows that $\operatorname{\mathbf{SCO}}$ strategies always exists in determined (well-formed as per their terminology) games, i.e., you can always find a strategy $\sigma$ such that $\operatorname{aVal}(h)=\operatorname{aVal}(h,\sigma)$ and $\operatorname{cVal}(h)=\operatorname{cVal}(h,\sigma)$ for every $h\in\operatorname{Plays}^{v}(\sigma)$.
@@ Appendix B Note on $\operatorname{\mathbf{SCO}}$ vs $\operatorname{\mathbf{SC}}$
Appendix C More admissible strategy properties
Lemma 11.
Given $\mathcal{G}$, and strategy $\sigma$ that is admissible, $\forall h\in\operatorname{Plays}^{h}(\sigma)$ such that $\operatorname{aVal}(h)<\infty$, $\sigma\not\equiv\sigma_{win}$
Proof.
Proof by Lemma 1. We first observe that an optimal winning strategy $\sigma_{win}$ satisfies the condition $\operatorname{aVal}(h,\sigma_{win})=\operatorname{aVal}(h)$. Thus, $\sigma_{win}$ enforce reaching a goal state for all Env player strategies $\tau\in\mathrm{T}$. Thus, every $\sigma_{win}$ ensures that every state for every play in $\operatorname{Plays}^{v}(\sigma_{win})$ belongs to $V_{win}$. Hence, from Lemma 1, we have that not all admissible strategies are $\operatorname{\mathbf{WCO}}$.
In Fig. 2, a Sys player strategy that commits to $v_{1}$ from $v_{0}$ is $\operatorname{\mathbf{WCO}}$. Thus, actions $v_{4}\to v_{5}$ and $v_{4}\to v_{7}$ belong to strategy that is $\operatorname{\mathbf{WCO}}$. But, only action $v_{4}\to v_{7}$ belongs to an admissible strategy as there exists a better payoff while enforcing the worst-case optimal payoff. Therefore, not every $\operatorname{\mathbf{WCO}}$ strategy is an admissible strategy. ∎
Thus, from Lemma 11, we have that admissible strategies are not equivalent to winning strategies.
C.1 When are $\operatorname{\mathbf{WCO}}$ admissible?
Another interesting characterization of admissible strategy is as follows. Given history $h$, if $\operatorname{acVal}(h)=\operatorname{aVal}(h)$, then the adversarial-cooperative value and adversarial value for $h$ are the same. This implies, strategy $\sigma$ that is $\operatorname{\mathbf{WCO}}$ is also $\operatorname{\mathbf{WCo-Op}}$. Thus, $\sigma$ that is worst-case optimal is admissible.
Lemma 12.
All worst-case optimal $(\operatorname{\mathbf{WCO}})$ strategies are admissible if $\operatorname{acVal}(h)=\operatorname{aVal}(h)$,
Proof.
A strategy $\sigma$ that witnesses $\operatorname{acVal}(h)$ is maximal in order. There does not exist another strategy that ensures a lower payoff while being $\operatorname{\mathbf{WCO}}$. There might exist $\sigma^{\prime}$ that achieves lower payoff against some adversary but will be receive higher payoff against another adversary, i.e., $\operatorname{aVal}(h,\sigma^{\prime})>\operatorname{aVal}(h)$. Hence, $\sigma^{\prime}$ does not dominate $\sigma$. Thus, if $\operatorname{acVal}(h)=\operatorname{aVal}(h)$, then all $\operatorname{\mathbf{WCO}}$ strategies are admissible. Further, if $\operatorname{acVal}(h)<\operatorname{aVal}(h)$, then a worst-case optimal strategy $\sigma$ is admissible if, and only if, $\operatorname{cVal}(h,\sigma)<\operatorname{aVal}(h)$ and $\operatorname{cVal}(h,\sigma)=~{}\operatorname{acVal}(h)$. ∎
Appendix D Clarification - Qualitative vs. Quantitative
D.1 Existing Qualitative Dominance Definition
These notions are common in the Formal Methods community where the objective is Boolean and the corresponding payoff is binary (win or lose (Berwanger 2007)).
Definition 14 (Dominance (Quant - very weakly dominating) (Aminof et al. 2021)).
Given two Sys strategies $\sigma,\sigma^{\prime}\in\Sigma$, we say $\sigma$ dominates $\sigma^{\prime}$, denoted by $\sigma\succeq\sigma^{\prime}$, if for every env strategy $\tau\in\mathrm{T}$, $P^{v_{0}}(\sigma^{\prime},\tau)\models\varphi\to P^{v_{0}}(\sigma,\tau)\models\varphi$.
Intuitively, given two strategies $\sigma$ and $\sigma^{\prime}$ for the agent, we say that $\sigma$ dominates $\sigma^{\prime}$ if $\sigma$ is as good $\sigma^{\prime}$ for some environment strategies $\tau$. We note that Strict Dominance was NOT defined using the definition env. (Aminof et al. 2021; Aminof, De Giacomo, and Rubin 2023).
Definition 15 (Strict Dominance (Quant - weakly dominating)).
Given two Sys strategies $\sigma,\sigma^{\prime}\in\Sigma$, we say $\sigma$ strictly dominates $\sigma^{\prime}$, denoted by $\sigma\succ\sigma^{\prime}$, if $\sigma\succeq\sigma^{\prime}$ and $\neg(\sigma^{\prime}\succeq\sigma)$.
We say $\sigma$ strictly dominates $\sigma^{\prime}$ if it does as good as $\sigma^{\prime}$ and is better than $\sigma^{\prime}$ in at least one case. Note, better $\to$ reaching a goal state for Two-player Reachability games.
Definition 16 (admissible (Berwanger 2007)).
A strategy is called admissible for the Sys player if it is not dominated ($\succ$; as per Def. 15) by any other strategy of the Sys player.
Definition 17 (Best-Effort (Aminof et al. 2021; Aminof, De Giacomo, and Rubin 2023; Faella 2009)).
A strategy is called Best-Effort for the Sys player if it is not dominated ($\succ$; as per Def. 15) by any other strategy of the Sys player.
So, by definition, in Qualitative settings, both Best-Effort and Admissible strategies are equivalent (Aminof et al. 2021; Aminof, De Giacomo, and Rubin 2023). The same is not true in the Quantitative settings.
Qualitative Quantitative
Dominance (Def. 14) Very weakly dominating (Def. 5)
strict dominance (Def. 15) ) weakly dominating (Def. 5)
N.A. strictly dominating (Def. 5)
Admissible $\equiv$ BE Admissible (BE - N.A.)
Table 3: Equivalent definition of dominance across qualitative (Aminof et al. 2021; Berwanger 2007) and quantitative papers (Leyton-Brown and Shoham 2008; Brenguier et al. 2016)

### v2

@@ B.17 Proof of Theorem 8
As $\operatorname{sVal}(v)=1$ only in the winning region, we have the following equivalent statement $\neg(v\in V_{win})\vee(v^{\prime}\in V_{win})$ where $v:=\operatorname{last}(h)$ and $v^{\prime}:=\delta(\operatorname{last}(h),\sigma(h))$. Thus, modifing the admissibility checking criteria in Line 9 of Alg. 1 to $\operatorname{cVal}(v^{\prime})<\min\{\operatorname{aValues}\}$ and $\neg(v\in V_{win})\vee(v^{\prime}\in V_{win})$ is sufficient to capture the admissibility criteria for $\operatorname{\mathbf{mSC}}$. Thus, from Thms. 5 and 6, we have that Alg. 4 returns the set of all admissible winning strategies, is sound and correct. ∎
B.18 Note on Alg. 2
Alg. 2 is the Value Iteration algorithm for computing optimal $\operatorname{Val}$s and optimal strategy from every state in $\mathcal{G}$ Brihaye et al. [2017][Adapted and Modified from Algo. 1]. Line 2 and Line 2 correspond to computation when the Env player is playing adversarially (maximally) and cooperatively (minimally), respectively. When Env is playing adversarially, Alg. 2 returns optimal $\operatorname{aVal}$ and worst-case optimal strategy $\operatorname{\mathbf{WCO}}$. For $\operatorname{aVal}$ computation, $\sigma_{win}$ is the optimal winning strategy from every state in $V_{win}$.
@@
B.19 Note on Alg. 3 and Alg. 4
Alg. 3 and Alg. 4 compute admissible and admissible winning strategies, respectively. The key components are unrolling the graph, running value iteration, and an algorithm for Depth-First Search (DFS). The DFS algorithm searches over $\mathcal{G}^{\prime}$ iteratively rather than recursively in order to avoid Python’s low recursion limit. Specifically, the algorithm explores states in preorder fashion. Alg. 2 gives the outline to compute optimal $\operatorname{aVal}$ and $\operatorname{cVal}$ and the corresponding witness strategies. See B.18 for more details. After we unroll the graph, we compute $\operatorname{aVal}$, $\operatorname{cVal}$, and $\operatorname{acVal}$ for every state and finally run the DFS algorithm for strategy synthesis.
@@
Appendix C Note on $\operatorname{\mathbf{SCO}}$ vs $\operatorname{\mathbf{SC}}$
Brenguier et al. [2016] define strongly-cooperative optimal strategies in order to show that every $\operatorname{\mathbf{SCO}}$ is admissible. Then, they proceed to shows that $\operatorname{\mathbf{SCO}}$ strategies always exists in determined (well-formed as per their terminology) games, i.e., you can always find a strategy $\sigma$ such that $\operatorname{aVal}(h)=\operatorname{aVal}(h,\sigma)$ and $\operatorname{cVal}(h)=\operatorname{cVal}(h,\sigma)$ for every $h\in\operatorname{Plays}^{v}(\sigma)$.